Consider the following measures based on independently drawn samples from normally distributed populations: (You may find it useful to reference the appropriate table: chi-square table or F table)

Sample 1: s21s12 = 221, and n1 = 16
Sample 2: s22s22 = 208, and n2 = 11

a. Construct the 95% interval estimate for the ratio of the population variances. (Round "F" value and final answers to 2 decimal places.)

Confidence interval _______ to _______

B. Using the confidence interval from Part (a), test if the ratio of the population variances differs from 1 at the 5% significance level.

The 95% confidence interval the value 1. Thus, we
conclude that the population variances differ at the 5% significance level.
Answer #1

a) The 95% confidence interval is given by

Fa/2,n1-1,n2-1 s2 2

Here c = 0.95, alpha =1-c=1-0.95=0.05

F/2,n2-1,n1-1 0.05/2,10,15 0.025,10,15 3.00

Fa/2n1-1,n2-1Fo.05/2,15,10Fo.025,15,10-3.52

Hence the required confidence interval is

<3.06.208 3,06 * 221 208

0.28 < 1 < 3.25

B. The hypothesis are

H0:rac{sigma _{1}^{2}}{sigma _{2}^{2}}=1: : : : : : v/s: : : : : : : : H1:rac{sigma _{1}^{2}}{sigma _{2}^{2}} eq 1

If 1 lies in the above interval we say that population variances does not differ from 1.

Here interval is ( 0.28, 3.25 ) . 1 does not lie in the interval. Hence we conclude that population variances differs from 1, at 5% level of significance.